Please help me my friends all have the same answer but I don't.
7.)How many quarts of a 60% alcohol solution must be added to 40 quarts of a 20% alcohol solution to obtain a mixture whic is 30% alcohol? All percentages are by volume.
6.) What weight of water must be evaporated from 40lb of a 20% salt solution to produce a 50% solution? All percentages are by weight.
thank you

Answer by checkley77(12844)

7) .6x+40*.2=.3(40+x)
.6x+8=12+.3x
.6x-.3x=12-8
.3x=4
x=4/.3
x=13.33 quarts of 60% alcohol is needed.
proof:
.6*13.33+8=.3(40+13.33)
8+8=.3*53.33
16=16
---------------------------------------------------
6) 40*.2=8 pounds of salt.
For a 50% solution you need 2*8=16 pounds of water.
40-16=24 pounds of water must be evaporated.